OpenStudy (anonymous):

(Function)-- Find f(x)*g(x) if f(x)=8x+1 and g(x)= (-5x)+9

OpenStudy (anonymous):

f(x)*g(x) = (8x+ 1) (-5x+9) here we need to use distribute property , (a+ b) (c+d) = a*c+a*d+b*c+b*d